OpenStudy (anonymous):

(8v+6)/(5) + (7v-8)/(4) = 26

OpenStudy (anonymous):

v = ?

OpenStudy (anonymous):

Multiply by 20 both sides

OpenStudy (mertsj):

Multiply every term by the common denominator which is 20

OpenStudy (anonymous):

@sudonick

OpenStudy (anonymous):

(8v+6) 5 + (7v−8) 4 =26 I will try to solve your equation. 3.35v−0.8=26(Simplify both sides of the equation) 3.35v−0.8+0.7999999999999998=26+0.7999999999999998(Add 0.7999999999999998 to both sides) 3.35v=26.8 (3.35v) 3.35 = 26.8 3.35 (Divide both sides by 3.35) v=8

OpenStudy (anonymous):

v indeed equals 8

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

ha - knew it